The amount of power dissipated in a 10 ohm resistor with 1 amp and 10 volts is _____________.
Answer #1

We know power dissipated by a resistor is

P = IV

P= I^(2)R

Here I = 1A, R=10ohm, and V=10volts

So, the power dissipated will be from above relations and given values:

P = 1A×10V = 10W

Or P = 1^(2)×10 = 10W